OpenStudy (hhopke.student):

College Algebra/Alg. 2 Review- operations with rational expressions (polynomials)

OpenStudy (hhopke.student):

\[\frac{ (-3mn)^2*64(m^2n^3)^2 }{ 16m^2n^4(mn^2)^3 } * \frac{ (3m^2n^3)^2 }{ 24(m^2n^2)^4 }\]

OpenStudy (hhopke.student):

The answer is 27/2mn^7 but I don't know how to get there.

TheSmartOne (thesmartone):

first use these properties to simplify \(\Large (a^b)^c = a^{bc}\) \(\Large a^b\times a^c = a^{b+c}\)

OpenStudy (hhopke.student):

\[\frac{ 9m^2n^2*64m^6n^3*9m^4n^6 }{ 16m^2n^4*m^3n^6*24m^8n^8 }\]

OpenStudy (hhopke.student):

\[\frac{ 81m^6n^8*64m^6n^3 }{ (16m^5(n^10))*24m^8n^8 }\]

OpenStudy (hhopke.student):

now I have \[\frac{ 27*4m }{ n^7 }\]

TheSmartOne (thesmartone):

you made a mistake right here http://prntscr.com/c6a6en it should be (m^2n^3)^2 = m^4n^6

TheSmartOne (thesmartone):

\[\frac{ 9m^2n^2*64m^4n^6*9m^4n^6 }{ 16m^2n^4*m^3n^6*24m^8n^8 }\] simplify it from here now :)

OpenStudy (hhopke.student):

ookay. Thanks!

TheSmartOne (thesmartone):

first simplify the numerator using first use these properties to simplify \(\Large a^b\times a^c = a^{b+c}\) and then the denominator and then use this formula \(\Large\frac{a^b}{a^c} = a^{b-c}\)